A store owner want to develop a new snack mix by mixing chocolate and trail mix. How many pounds of chocolate costing $14.50 per pound should be mixed with trail mix costing $4.40 per pound to create a 40 pound mixture worth $8.19 per pound. (round to the nearest pound)

Answer by richwmiller(17219)

40*$8.19=327.60
We know the total amount of the candy . 40
x+y=40
We know that the total cost for the candy is 327.6
14.50*x+4.40*y= 327.6
x=40-y
We substitute for x
14.50*(40-y)+4.40*y= 327.6
We multiply out
580-14.50*y+4.40*y= 327.60
We combine like terms.
-10.1*y=327.60-580
-10.1*y=-252.40
-10.1*y=-252.40
Isolate y
y=-252.4/-10.10
y=24.990099009901=25 lbs chocolate candy at $ 4.40 $ is-worth $109.956435643564
Calculate x
x=40-24.990099009901
x=15.009900990099=15 lbs trail candy at $ 14.50 $ is-worth $217.643564356436
We check
14.50*15.009900990099+4.40*24.990099009901=327.60
217.643564356436+109.956435643564=327.60
327.6=327.60
Since 327.60 = 327.60 is TRUE, all is well